OpenStudy (anonymous):

How do I set this problem up? 1) How many even 3-digit positive integers can be written using the digits1, 2, 4, 7 & 8 ?

OpenStudy (cwrw238):

are duplicates allowed?

OpenStudy (cwrw238):

eg 224 , 444 ?

OpenStudy (cwrw238):

if not then proceed as follows: the numbers must end in 2 , 4 or 8 for each of these numbers we have a number of permutations of 2 from the other 4 numbers - that is 4P2 = 12

OpenStudy (cwrw238):

can u continue ?

OpenStudy (anonymous):

why did u set it equal to 12?

OpenStudy (cwrw238):

the number of permutaions of 2 from 4 = 4! / 2! = 24 /2 = 12

OpenStudy (cwrw238):

and has there are 3 even nunbers the number of possibilities = 3 * 12 - this is true if no duplicates or triplicates are allowed

OpenStudy (anonymous):

im confused.. the answer is 75 so im not sure how to get it